# Solving the Optimal Experiment Design Problem with Mixed-Integer Convex Methods

### v5

@@ 2.1.1 D-Criterion
as the D-Optimal Experiment Design Problem (D-Opt). Observe that the objective is equivalent to minimizing the determinant of the dispersion matrix. This determinant is also called the generalized variance of the parameter $\bm{\theta}$ (Pukelsheim, 2006). A maximal value of $\det X$ corresponds to a minimal volume of standard ellipsoidal confidence region of $\bm{\theta}$ (Ponte et al., 2023). Additionally, the D-criterion is invariant under reparameterization, see Pukelsheim (2006).
2.1.2 A-Criterion
@@
For parameters with a physical interpretation, the A-optimality criterion is a good choice as it amounts to minimizing the average of the variances of $\bm{\theta}$ (Pukelsheim, 2006). Using the $\log$ rules for the objective, the A-Optimal Experiment Design Problem (A-Opt) can be stated as
@@
$\begin{split}\min\limits_{\mathbf{x}}\,&\operatorname{Tr}\left((X(\mathbf{x}))^{-1}\right)\\ \text{s.t. }\,&\,\mathbf{x}\in\mathcal{P}\cap{\mathbb{Z}}^{m}.\end{split}$ (A-Opt)
Note that the D-Optimal Problem is known to be $\mathcal{NP}$-hard since the eighties (Welch, 1982). $\mathcal{NP}$-hardness of the A-Optimal Problem and the D-Fusion Problem was proved only recently, see (Nikolov et al., 2022) and (Ponte et al., 2023), respectively. The hardness of the A-Fusion Problem is still open though the authors conjecture that it too is $\mathcal{NP}$-hard.
3 Convergence guarantees for the continuous subproblems
To apply the new framework Boscia, we need to guarantee that the Frank-Wolfe algorithm converges on the continuous subproblems. The conventional property guaranteeing convergence is $L$-smoothness of the objective, that is
@@ 3 Convergence guarantees for the continuous subproblems
$\exists L\in{\mathbb{R}}_{>0}\forall\mathbf{x},\mathbf{y}\in\mathcal{P}\text{ s.t. }\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|\leq L\left\|\mathbf{x}-\mathbf{y}\right\|.$
In the case of both the D-Optimal Problem and the A-Optimal Problem, not every point $\mathbf{x}\in\mathcal{P}$ is domain-feasible for the objective functions due to the corresponding information matrix being singular. For such points $\mathbf{x}$, we define $-\log\det(X(\mathbf{x}))=\infty$ and $\operatorname{Tr}\left((X(\mathbf{x}))^{-1}\right)=\infty$ for the D-criterion and A-criterion, respectively. Thus, the objective functions of (A-Opt) and (D-Opt) are not $L$-smooth over the feasible region.
@@
Lacking $L$-smoothness on the feasible region, we require a different property that guarantees convergence of the Frank-Wolfe algorithm for the Optimal Problems under both criteria. By (Carderera et al., 2024), the Frank-Wolfe algorithm also converges (with similar convergence rates) if the objective is generalized self-concordant.
Definition 4 (Generalized Self-Concordance Sun & Tran-Dinh (2019))
@@
$|f^{\prime\prime\prime}(x)|\leq M_{f}f^{\prime\prime}(x)^{\frac{\nu}{2}}.$
@@
By (Sun & Tran-Dinh, 2019, Proposition 2), the composition of a generalized self-concordant function with a linear map is still generalized self-concordant. Hence, it suffices that we show that the functions $f(X)=-\log\det(X)$ and $g(X)=\operatorname{Tr}\left(X^{-p}\right)$, $p>0$, are generalized self-concordant for $X$ in the PD cone. Note that $f$ is the logarithmic barrier for the PSD cone which is known to be self-concordant (Nesterov & Nemirovskii, 1994). Convergence of the Frank-Wolfe algorithm for the (D-Opt) Problem is therefore guaranteed.
For the (A-Opt) Problem, we can show that $g$ is self-concordant on a part of the PD cone, namely one characterized by an upper bound on the maximum eigenvalue.
Theorem 3.1

### v6

@@ 2.1 The A-, D- and GTI-optimal experiment design problem
Observe that it is equivalent to minimize $\log\det(X(\mathbf{x})^{-1})=\log\det D(\mathbf{x})$, so the determinant of the dispersion matrix. This is also called the generalized variance of the parameter $\bm{\theta}$ (Pukelsheim, 2006). Thus, a maximal value of $\det X$ corresponds to a maximal volume of standard ellipsoidal confidence region of $\bm{\theta}$ (Ponte et al., 2025). Additionally, the D-criterion is invariant under reparameterization (Pukelsheim, 2006).
A-Criterion
@@
For parameters with a physical interpretation, the A-optimality criterion is a good choice (Pukelsheim, 2006) as it amounts to minimizing the average of the variances of $\bm{\theta}$. Geometrically, the A-criterion amounts to minimizing the diagonal of the bounding box of the ellipsoidal confidence region (Sagnol, 2010). The resulting optimization problem can be stated in two different ways, keeping the $\log$ and using the $\log$ rules and the positive homogeneity of the information functions.
@@
$\begin{split}\min\limits_{\mathbf{x}}\,&\log\left(\operatorname{Tr}\left((X(\mathbf{x}))^{-1}\right)\right)\\ \text{s.t.}\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{N}}_{0}^{m}\end{split}$ (logA-Opt)
$\begin{split}\min\limits_{\mathbf{x}}\,&\operatorname{Tr}\left((X(\mathbf{x}))^{-1}\right)\\ \text{s.t.}\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{N}}_{0}^{m}\end{split}$ (A-Opt)
We test both formulation within the newly proposed framework. Both formulations can be generalized for a real number $p>0$:
$\begin{split}\min\limits_{\mathbf{x}}\,&\log\left(\operatorname{Tr}\left((X(\mathbf{x}))^{-p}\right)\right)\\ \text{s.t.}\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{N}}_{0}^{m}\end{split}$ (logGTI-Opt)
$\begin{split}\min\limits_{\mathbf{x}}\,&\operatorname{Tr}\left((X(\mathbf{x}))^{-p}\right)\\ \text{s.t.}\,&\sum_{i=1}^{m}x_{i}=N\\ &\mathbf{l}\leq\mathbf{x}\leq\mathbf{u}\\ &\mathbf{x}\in{\mathbb{N}}_{0}^{m}\end{split}$ (GTI-Opt)
We call the corresponding problem the Generalized-Trace-Inverse Optimal Experiment Design Problem, short GTI-Optimal Problem.
Remark.
We conjecture that OEDP under the GTI-criterion is $\mathcal{NP}$-hard for any $p>0$.
Note that the objectives of (D-Opt), (A-Opt), (logA-Opt), (logGTI-Opt) and (GTI-Opt) are only well defined if the information matrix $X(\mathbf{x})$ has full rank.
Remark.
The Fusion Problems are created by replacing $X(\mathbf{x})$ with $X_{C}(\mathbf{x})$ in (D-Opt), (A-Opt), (logA-Opt), (logGTI-Opt) and (GTI-Opt), respectively.
For convenience in proofs in the next section, we define
@@
$\mathcal{P}:=[\mathbf{l},\mathbf{u}]\cap\left\{\mathbf{x}\in{\mathbb{R}}^{m}_{\geq 0}\,\middle|\,\sum x_{i}=N\right\}$ (Convex Hull)
as the feasible region of the OEDPs without the integer constraints, i.e. the convex hull of all feasible integer points. Thus, the feasible region with integer constraints will be noted by $\mathcal{P}\cap{\mathbb{Z}}$.
We denote the domain of the objective functions by
$D:=\left\{\mathbf{x}\in{\mathbb{R}}^{m}_{\geq 0}\,\middle|\,X(\mathbf{x})\in\mathbb{S}_{++}^{n}\right\}$ (Domain)
Throughout the rest of this paper, we denote by
$f(X)=-\log\det(X)$
$g(X)=\operatorname{Tr}\left(X^{-p}\right)$
$k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$
where $X\in\mathbb{S}_{++}^{n}$ and $p>0$. With abusive of notation, we use the same identifiers for the objectives of the optimization problems.
$f(\mathbf{x})=-\log\det(X(\mathbf{x}))$
$g(\mathbf{x})=\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)$
$k(\mathbf{x})=\log\left(\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)\right)$
3 Properties of the problems and convergence guarantees
To ensure that the newly proposed framework in Hendrych et al. (2025b) is applicable to the previously introduced problems, we have to show that the Frank-Wolfe algorithms used as the subproblem solvers converge on the continuous subproblems. Frank-Wolfe normally requires the objective function to be convex and $L$-smooth. The convexity part is trivial, for proofs see Section A.2. The $L$-smoothness property only holds for the Fusion Problems on the whole feasible region. For the Optimal Problems, we will show local $L$-smoothness in Section 3.1 and an alternative property guaranteeing convergence of Frank-Wolfe, namely (generalized) self-concordance, in Section 3.2.
Lastly, we establish sharpness or the Hölder error bound condition of the objective functions, therefore yielding improved convergence rates for Frank-Wolfe methods compared to the standard $\mathcal{O}(1/\epsilon)$ rate where $\epsilon$ is the additive error in the primal gap, see Section 3.3.
3.1 Lipschitz smoothness
Definition 3.3.
A function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is called $L$-smooth if its gradient $\nabla f$ is Lipschitz continuous. That is, there exists a finite $L\in{\mathbb{R}}_{>0}$ with
$\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|_{2}\leq L\left\|\mathbf{x}-\mathbf{y}\right\|_{2}$
Since the functions of interest are in $\mathcal{C}^{3}$, we can use an alternative condition on $L$-smoothness:
$LI\succcurlyeq\nabla_{xx}f(\mathbf{x}),$
i.e. the Lipschitz constant $L$ is an upper bound on the maximum eigenvalue of the Hessian.
In Section A.3, the reader can find $L$-smoothness proofs for $f(X)$, $g(X)$ and $k(X)$ over
$\{X\in\mathbb{S}_{++}^{n}\,\mid\,\delta\geq\lambda_{\min}(X)\}.$
The lemmas are not immediately useful for us, in general we cannot bound the minimum eigenvalue of the information matrix. In fact, we will see that we have $L$-smoothness on the whole convex feasible region $\mathcal{P}$ only for the Fusion Problems. Due to the positive definiteness of the matrix $C$, the information matrix in the Fusion case, $X_{C}(\mathbf{x})$, is always bounded away from the boundary of the PD cone and we do find a lower bound on the minimum eigenvalue of $X_{C}(\mathbf{x})$. For Optimal Problems, $X(\mathbf{x})$ can be singular for some $\mathbf{x}\in\mathcal{P}$, i.e. the minimum eigenvalue of $X(\mathbf{x})$ cannot be bounded away from 0 on $\mathcal{P}$. However, we can show $L$-smoothness in a local region for the Optimal Problems.
We start with the $L$-smoothness proofs for the objectives of the Fusion Problems.
Theorem 3.4.
The functions $f_{F}(\mathbf{x}):=-\log\det\left(X_{C}(\mathbf{x})\right)$ and $g_{F}(\mathbf{x}):=\operatorname{Tr}\left(\left(X_{C}(\mathbf{x})\right)^{-p}\right)$ for $p\in{\mathbb{R}}_{>0}$ are $L$-smooth on $x\in{\mathbb{R}}_{\geq 0}$ with Lipschitz constants
$L_{f_{F}}=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}}{\lambda_{\min}(C)^{2}}$
and
$L_{g_{F}}=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}},$
respectively.
Proof.
For better readability and simplicity, we ignore the dependency of $X$ on $\mathbf{x}$ for now. As stated above, we can prove $L$-smoothness by showing that the largest eigenvalue of the Hessians is bounded for any $\mathbf{x}$ in the domain. We have
$\nabla^{2}f_{F}(\mathbf{x})$ $=(AA^{\intercal})\circ(AX_{C}^{-2}A^{\intercal})$
and
$\nabla^{2}g_{F}(\mathbf{x})$ $=p(p+1)\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right).$
Observe that the Hessian expressions are very similar. Hence, we will only present the proof for $g_{F}$. The proof of $f_{F}$ follows the same argumentation. By Horn (1990, page 95) it is known that the maximum eigenvalue of a Hadamard product between two positive semi-definite matrices $N$ and $M$, $\lambda_{\max}(N\circ M)$, is upper bounded by the maximum diagonal entry of $N$ and the largest eigenvalue of $M$, so
$\lambda_{\max}(N\circ M)$ $\leq\max_{i}N_{ii}\lambda_{\max}(M).$ (8)
In our case, $N=AA^{\intercal}$ and $M=AX_{C}^{-p-2}A^{\intercal}$. By Corollary A.19 in Section A.1, we have that $X_{C}^{-p-2}$ is positive definite. Thus, both factors are semi-definite and we may use the provided bound.
$\lambda_{\max}\left(\nabla^{2}g_{F}(\mathbf{x})\right)$ $\leq\max\limits_{i}(AA^{\intercal})_{ii}\lambda_{\max}(AX_{C}^{-p-2}A^{\intercal})$
Note that the first factor is simply
For the second factor, we can exploit positive semi-definiteness with the spectral norm.
$\leq\left\|A\right\|_{2}^{2}\left\|X_{C}^{-p-2}\right\|_{2}$
By Lemma A.18, we have $\lambda(A^{p})=\lambda(A)^{p}$.
$=\left\|A\right\|_{2}^{2}\lambda_{\max}(X_{C}^{-1})^{2+p}$
$=\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(X_{C})^{2+p}}$
By the definition of $X$ and by Weyl’s inequality on the eigenvalues of sums of positive semi-definite matrices Weyl (1912), we have that
Hence,
Combining (8) with (9) and (10) yields
$\lambda_{\max}\left(\nabla^{2}g_{F}(\mathbf{x})\right)$ $\leq p(p+1)\left(\max\limits_{i}v_{i}^{\intercal}v_{i}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}}.$ (11)
For the $\log$ variant of the A-criterion, we can also bound the maximum eigenvalue of the hessian $\nabla^{2}k_{F}$.
Theorem 3.5.
The function $k_{F}(\mathbf{x}):=\log\operatorname{Tr}\left(\left(X_{C}(\mathbf{x})\right)^{-p}\right)$ is $L$-smooth on $\mathcal{P}$ with Lipschitz constant
$L_{k_{F}}=\frac{a_{C}^{p}}{n}p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(C)^{2+p}}$
where $a_{C}$ is the upper bound on the maximum eigenvalue of $X_{C}(\mathbf{x})$, that is
$a_{C}=\lambda_{\max}(C)+\max_{i\in[m]}u_{i}\max_{j\in[m]}\left\|v_{j}\right\|_{2}^{2}$
with $u_{i}$ denoting the upper bound of variable $x_{i}$.
Proof.
The Hessian of $k_{F}$ is given by
$\nabla^{2}k_{F}$ $=p\frac{(p+1)\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right)\operatorname{Tr}\left(X_{C}^{-p}\right)-p\operatorname{diag}\left(AX_{C}^{-p-1}A^{\intercal}\right)\operatorname{diag}\left(AX_{C}^{-p-1}A^{\intercal}\right)^{\intercal}}{\operatorname{Tr}\left(X_{C}^{-p}\right)^{2}}.$
As before, we omit the dependency on $\mathbf{x}$ for better readability. Note that both parts of the nominator are symmetric matrices. Thus, we can use the upper bound on the eigenvalues of sums of symmetric matrices in Lemma A.28. Further, note that for any $\mathbf{z}\in{\mathbb{R}}^{n}$ the matrix $\mathbf{z}\mathbf{z}^{\intercal}$ is PSD and its maximum eigenvalue is $\left\langle\mathbf{z},\mathbf{z}\right\rangle$.
$\lambda_{\max}\left(\nabla^{2}k_{F}\right)$ $\leq\frac{p(p+1)}{\operatorname{Tr}\left(X_{C}^{-p}\right)}\lambda_{\max}\left(\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right)\right)$
An upper bound on the maximum eigenvalue of $\left(AA^{\intercal}\right)\circ\left(AX_{C}^{-p-2}A^{\intercal}\right)$ is already provided by Theorem 3.4. Observe that
$\operatorname{Tr}\left(X_{C}^{-p}\right)\geq n\lambda_{\min}\left(X_{C}^{-p}\right)=\frac{n}{\lambda_{\max}(X_{C})^{p}}\geq\frac{n}{a_{C}}.$
Assembling everything yields the desired bound. ∎
As stated before, the proofs of Theorems 3.4 and 3.5 work because we can bound the minimum eigenvalue of the Fusion information matrix $X_{C}$ away from zero thanks to the regular matrix $C$. This is not the case of the information matrices $X(\mathbf{x})$ of the Optimal Problems as they can be singular for some $x\in\mathcal{P}$. Therefore, the objectives of the Optimal Problems are not $L$-smooth over $\mathcal{P}$. We can show, however, that the objectives are $L$-smooth on a restricted area of the feasible region.
Remember that $D$ denotes the domain of the objectives, see (Domain). Since the Frank-Wolfe algorithms make monotonic progress, it is helpful to show that given an initial point $\mathbf{x}_{0}$ in the domain $D$ the minimum eigenvalue is bounded in the region
$\mathcal{L}_{0}=\left\{\mathbf{x}\in D\cap\sum\limits_{i=1}^{m}x_{i}=N\,\middle|\,(*)(\mathbf{x})\leq(*)(\mathbf{x}_{0})\right\}$
where ()(*)( ∗ ) is a placeholder for $f(\mathbf{x})$, $g(\mathbf{x})$ and $k(\mathbf{x})$, respectively. Using the bound provided by Merikoski & Virtanen (1997, Theorem 1) and the fact that the objective value is decreasing, we can bound the minimum eigenvalue by quantities depending on the experiment matrix $A$ and the start point $\mathbf{x}_{0}$. Let further $X_{0}=X(\mathbf{x}_{0})$.
Theorem 3.6.
The functions $f(\mathbf{x}):=-\log\det\left(X(\mathbf{x})\right)$, $g(\mathbf{x}):=\operatorname{Tr}\left(\left(X(\mathbf{x})\right)^{-p}\right)$ and $k(\mathbf{x}):=\operatorname{Tr}\left(\left(X(\mathbf{x})\right)^{-p}\right)$ for $p>0$ are $L$-smooth locally on their respective $\mathcal{L}_{0}$ for a given initial point $x_{0}$ with
$L_{f}=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}\left(mN\max_{i}\left\|v_{i}\right\|_{2}^{2}\right)^{n-1}}{\left(n-1\right)^{n-1}\det X_{0}^{2}}$
$L_{g}=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\left\|A\right\|_{2}^{2}\left(\operatorname{Tr}\left(X_{0}^{-1}\right)\right)^{2+p}$
and
$L_{k}=\frac{a^{p}}{n}p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\left\|A\right\|_{2}^{2}\left(\operatorname{Tr}\left(X_{0}^{-1}\right)\right)^{2+p}$
where $a=\max_{i\in[m]}u_{i}\max_{j\in[m]}\left\|v_{j}\right\|_{2}^{2}$ is a upper bound on the maximum eigenvalue of $X(\mathbf{x})$.
Proof.
As seen in the previous proofs, we have
$L_{f}$ $=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}}{\lambda_{\min}(X(\mathbf{x}))^{2}}$
$L_{g}$ $=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(X(\mathbf{x}))^{2+p}}.$
and
$L_{k}$ $=\frac{a^{p}}{n}p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\left\|A\right\|_{2}^{2}}{\lambda_{\min}(X(\mathbf{x}))^{2+p}}$
To bound the Lipschitz constant, we need lower bound the minimum eigenvalue of $X(\mathbf{x})$. We start with $L_{f}$ and find by using the bound provided by Merikoski & Virtanen (1997, Theorem 1)
$\lambda_{\min}(X)$ $\geq\left(\frac{n-1}{\operatorname{Tr}(X)}\right)^{n-1}\det X.$
The trace can be upper bounded by $\operatorname{Tr}(X)\leq mN\max_{i}\left\|\mathbf{v}_{i}\right\|_{2}^{2}$ and by knowing that $-\log\det(X(\mathbf{x}))\leq-\log\det(X_{0})$, we find
$\lambda_{\min}(X)$ $\geq\left(\frac{n-1}{mN\max_{i}\left\|\mathbf{v}_{i}\right\|_{2}^{2}}\right)^{n-1}\det X_{0}.$
Thus,
$L_{f}$ $=\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\frac{\|A\|_{2}^{2}\left(mN\max_{i}\left\|\mathbf{v}_{i}\right\|_{2}^{2}\right)^{n-1}}{\left(n-1\right)^{n-1}\det X_{0}^{2}}.$
To lowerbound the minimum eigenvalue of $X(\mathbf{x})$ w.r.t. the objectives $g(\mathbf{x})$ and $k(\mathbf{x})$, we first need
$\operatorname{Tr}(X^{-p-2})$ $=\sum\limits_{i=1}^{n}\frac{1}{\lambda_{i}(X)^{p+2}}=\sum\limits_{i=1}^{n}\frac{1}{\lambda_{i}(X)^{2}}\frac{1}{\lambda_{i}(X)^{p}}$
$\leq\frac{1}{\lambda_{\min}(X)^{2}}\operatorname{Tr}(X^{-p}).$
Note that the $\log$ is preserving ordering. Hence, we can use the fact that $\operatorname{Tr}(X^{-p})\leq\operatorname{Tr}(X_{0}^{-p})$ and we find
$\leq\frac{1}{\lambda_{\min}(X)^{2}}\operatorname{Tr}(X_{0}^{-p}).$
Exploiting positive definiteness and the spectral norm, we have
$\lambda_{\min}(X)^{p+2}$ $=\lambda_{\min}\left(X^{\frac{p+2}{2}}\right)^{2}=\frac{1}{\lambda_{\max}\left(X^{-\frac{p+2}{2}}\right)^{2}}$
$=\frac{1}{\left\|X^{-\frac{p+2}{2}}\right\|_{2}^{2}}\geq\frac{1}{\left\|X^{-\frac{p+2}{2}}\right\|_{F}^{2}}=\frac{1}{\operatorname{Tr}\left(X^{-p-2}\right)}$
$\lambda_{\min}(X)^{p+2}$ $\geq\frac{\lambda_{\min}(X)^{2}}{\operatorname{Tr}(X^{-p})}\geq\frac{\lambda_{\min}(X)^{2}}{\operatorname{Tr}\left(X_{0}^{-p}\right)}\geq\frac{\lambda_{\min}(X)^{2}}{\operatorname{Tr}\left(X_{0}^{-1}\right)^{p}}.$
Ultimately, we find
$\lambda_{\min}(X)$ $\geq\frac{1}{\operatorname{Tr}\left(X_{0}^{-1}\right)}$
and the smooth constants are
$L_{g}$ $=p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\left\|A\right\|_{2}^{2}\left(\operatorname{Tr}\left(X_{0}^{-1}\right)\right)^{2+p}$
and
$L_{k}$ $=\frac{a^{p}}{n}p(p+1)\left(\max\limits_{i}(AA^{\intercal})_{ii}\right)\left\|A\right\|_{2}^{2}\left(\operatorname{Tr}\left(X_{0}^{-1}\right)\right)^{2+p}.$
Thereby, we have local smoothness constants for the objectives $f$, $g$ and $k$. ∎
To summarize, we have shown $L$-smoothness of the objectives of the Fusion Problems which guarantees convergence of Frank-Wolfe. For the Optimal Problems, we have shown local $L$-smoothness on a restricted area of the feasible region. This, however, does not suffice to guarantee convergence of Frank-Wolfe. In the next section, we will show that the objectives satisfy an alternative condition that ensures convergence of Frank-Wolfe.
3.2 Generalized self-concordance
@@
By Carderera et al. (2021), the Frank-Wolfe algorithm also converges (with similar convergence rates to the $L$-smooth case) if the objective is generalized self-concordant.
Definition 3.7 (Generalized Self-Concordance Sun & Tran-Dinh (2019)).
@@
$|f^{\prime\prime\prime}(x)|\leq M_{f}\left(f^{\prime\prime}(x)\right)^{\frac{\nu}{2}}.$
@@
By Sun & Tran-Dinh (2019, Proposition 2), the composition of a generalized self-concordant function with a linear map is still generalized self-concordant. Hence, it suffices that we show that the functions $f(X)=-\log\det(X)$ and $g(X)=\operatorname{Tr}\left(X^{-p}\right)$ and $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$, $p>0$, are generalized self-concordant for $X$ in the PD cone.
Notice that $f$ is the logarithmic barrier for the PSD cone which is known to be self-concordant (Nesterov & Nemirovskii, 1994). Convergence of Frank-Wolfe for the (D-Opt) Problem is therefore guaranteed. For convergence for the (GTI-Opt) Problem and (A-Opt) Problem, we can show that $g$ is self-concordant on a part of the PD cone.
Theorem 3.8.